OpenStudy (anonymous):

x/2 + y/5 = 1 I need to find slope and intercepts of the line. How do I move the x/2 to the other side?

OpenStudy (anonymous):

Multiply both sides by 2 and 5. 5x + 2y = 1 2y = -5x + 1 y = -5/2x + 1/2

OpenStudy (anonymous):

so do I multiply the x/2 by 2 and the y/5 by 5?....I issue I am having is with the fractions.

OpenStudy (anonymous):

Oh no, this is what happens. x/2 + y/5 = 1 and then x(5)(2)/2 + y(5)(2)/5 = (5)(2) The 2's under the x cancel and the 5 for the y cancel. 5x + 2y = 10